# LUSIFER: Language Universal Space Integration for Enhanced Multilingual Embeddings with Large Language Models

### v1

@@ Appendix A Training Details
Table 5: Number of samples used in each dataset for training. The number of negative samples is included in the total number of samples.
Appendix B Detailed Results
In this section, we provide detailed results of LUSIFER and E5-Mistral on all benchmark datasets for each language.
Es Datasets E5-Mistral LUSIFER
AmazonReviewsClassification 42.69 50.41
MassiveIntentClassification 69.67 68.93
MassiveScenarioClassification 74.63 73.41
MTOPIntentClassification 72.16 80.13
MultilingualSentimentClassification 87.91 91.01
TweetSentimentClassification 49.73 58.55
SpanishNewsClassification 89.5 87.81
PawsXPairClassification 61.19 62.82
XNLI 77.34 60.49
SpanishNewsClusteringP2P 42.28 43.85
MLSUMClusteringP2P 47.54 44.36
MLSUMClusteringS2S 47.11 41.56
SIB200ClusteringS2S 31.01 44.42
MultiEURLEXMultilabelClassification 6.16 3.87
BelebeleRetrieval 83.92 81.4
MintakaRetrieval 48.77 18.17
STS17 87.18 80.84
STS22 71.79 70.66
STSBenchmarkMultilingualSTS 84.31 79.89
Avg. 61.84 60.14
Table 6: Detailed results of E5-Mistral and LUSIFER on the Spanish benchmark datasets.
En Datasets E5-Mistral LUSIFER
AmazonCounterfactualClassification 78.69 72.45
AmazonPolarityClassification 95.91 94.3
AmazonReviewsClassification 55.79 55.46
Banking77Classification 88.23 87.33
EmotionClassification 49.77 74
ImdbClassification 94.78 92.52
MassiveIntentClassification 80.57 75.64
MassiveScenarioClassification 82.39 78
MTOPDomainClassification 96.12 96.81
MTOPIntentClassification 86.11 87.34
ToxicConversationsClassification 69.59 82.84
TweetSentimentExtractionClassification 63.72 72.74
SprintDuplicateQuestions 95.66 90.99
TwitterSemEval2015 81.62 68.49
TwitterURLCorpus 87.75 85.35
ArxivClusteringP2P 50.45 35.6
ArxivClusteringS2S 45.5 22.25
BiorxivClusteringP2P 43.53 39.93
BiorxivClusteringS2S 40.24 29.3
MedrxivClusteringP2P 38.19 41.2
MedrxivClusteringS2S 37.45 35.53
RedditClustering 57.71 39.94
RedditClusteringP2P 66.49 53.4
StackExchangeClustering 73.1 46.41
StackExchangeClusteringP2P 45.91 39.7
TwentyNewsgroupsClustering 54.31 38.5
AskUbuntuDupQuestions 66.98 60.56
MindSmallReranking 32.6 24.55
SciDocsRR 86.33 34.94
StackOverflowDupQuestions 54.91 46.04
ArguAna 61.88 74.15
ClimateFEVER 38.4 29.24
CQADupstackTexRetrieval 42.97 23.22
DBPedia 48.9 17.98
FEVER 87.8 82.77
FiQA2018 56.62 14.91
HotpotQA 75.7 49.04
MSMARCO 43.1 56.43
NFCorpus 38.59 5.48
NQ 63.5 42.95
QuoraRetrieval 89.62 89.1
SCIDOCS 16.27 5.53
SciFact 76.41 66.09
Touche2020 26.39 6.33
TRECCOVID 87.33 18.22
STS12 79.65 74.26
STS13 88.43 84.2
STS14 84.54 77.5
STS15 90.42 84.95
STS16 87.68 82.21
STS17 91.75 81.67
STS22 67.28 71.25
BIOSSES 82.64 84.22
SICK-R 80.76 78
STSBenchmark 88.6 84.18
SummEval 31.4 32.36
Avg. 67.69 57.20
Table 7: Detailed results of E5-Mistral and LUSIFER on the English benchmark datasets.
Ru Datasets E5-Mistral LUSIFER
GeoreviewClassification 46.92 43.79
HeadlineClassification 76.52 79.26
InappropriatenessClassification 59.35 63.15
KinopoiskClassification 60.67 60.57
MassiveIntentClassification 72.06 71.29
MassiveScenarioClassification 76.64 74.49
RuReviewsClassification 64.10 67.40
RuSciBenchGRNTIClassification 60.19 59.51
RuSciBenchOECDClassification 46.30 46.41
GeoreviewClusteringP2P 69.87 59.20
RuSciBenchGRNTIClusteringP2P 52.96 55.00
RuSciBenchOECDClusteringP2P 46.54 49.95
TERRa 57.45 54.24
RiaNewsRetrieval 71.39 49.61
RuBQRetrieval 38.04 43.48
RuSTSBenchmarkSTS 81.79 78.20
STS22 61.32 61.44
Avg. 61.30 59.82
Table 8: Detailed results of E5-Mistral and LUSIFER on the Russian benchmark datasets.
Fr Datasets E5-Mistral LUSIFER
AmazonReviewsClassification 43.36 49.96
MTOPIntentClassification 70.39 79.14
MassiveIntentClassification 71.12 70.88
MassiveScenarioClassification 74.68 73.96
TweetSentimentClassification 50.23 62.62
SIB200Classification 72.45 79.51
FrenchBookReviews 46.77 48.07
PawsXPairClassification 62.15 65.93
RTE3 88.45 87.62
XNLI 76.60 62.75
MasakhaNEWSClusteringP2P 50.96 48.59
MasakhaNEWSClusteringS2S 52.08 63.12
MLSUMClusteringP2P 42.69 42.70
MLSUMClusteringS2S 42.60 41.51
HALClusteringS2S 24.21 24.16
SIB200ClusteringS2S 29.94 43.30
MultiEURLEXMultilabelClassification 5.00 3.51
BelebeleRetrieval 84.66 83.76
MintakaRetrieval 52.60 18.88
OpusparcusPC 94.58 90.63
STS17 84.66 82.19
SICKFr 79.12 74.22
STS22 76.50 73.77
STSBenchmarkMultilingualSTS 83.98 78.42
SummEvalFr 31.38 31.91
Avg. 59.65 59.24
Table 9: Detailed results of E5-Mistral and LUSIFER on the French benchmark datasets.
Vi Datasets E5-Mistral LUSIFER
MassiveIntentClassification 66.36 71.38
MassiveScenarioClassification 70.69 74.82
MultilingualSentimentClassification 69.30 81.30
SIB200Classification 70.20 78.58
VieStudentFeedbackClassification 73.02 77.39
XNLI 71.32 61.30
SIB200ClusteringS2S 32.93 46.79
BelebeleRetrieval 79.20 85.51
MLQARetrieval 32.43 54.61
VieQuADRetrieval 20.35 45.20
Avg. 58.58 67.69
Table 10: Detailed results of E5-Mistral and LUSIFER on the Vietnamese benchmark datasets.
Fa Datasets E5-Mistral LUSIFER
MassiveScenarioClassification 76.37 77.94
MassiveIntentClassification 71.98 73.32
MultilingualSentimentClassification 80.07 80.54
FarsTail 63.49 67.98
WikipediaRerankingMultilingual 75.60 78.75
WikipediaRetrievalMultilingual 67.77 78.49
Avg. 72.55 76.17
Table 11: Detailed results of E5-Mistral and LUSIFER on the Farsi benchmark datasets.
Id Datasets E5-Mistral LUSIFER
IndonesianMongabayConservationClassification 24.72 25.27
MassiveIntentClassification 69.51 71.38
MassiveScenarioClassification 72.89 74.62
SIB200Classification 80.88 80.44
indonli 50.00 50.22
SIB200ClusteringS2S 46.46 47.50
BelebeleRetrieval 81.10 87.56
SemRel24STS 40.40 40.57
Avg. 58.25 59.70
Table 12: Detailed results of E5-Mistral and LUSIFER on the Indonesian benchmark datasets.
Ar Datasets E5-Mistral LUSIFER
TweetEmotionClassification 53.74 49.03
ArEntail 77.63 84.15
XNLI 68.00 58.58
MintakaRetrieval 17.15 16.59
MLQARetrieval 28.32 47.90
STS17 75.13 71.44
STS22 61.01 61.54
Avg. 54.43 55.60
Table 13: Detailed results of E5-Mistral and LUSIFER on the Arabic benchmark datasets.
Fi Datasets E5-Mistral LUSIFER
FinToxicityClassification 53.78 62.23
MassiveIntentClassification 64.15 70.77
MassiveScenarioClassification 67.79 75.02
MultilingualSentimentClassification 72.42 83.59
SIB200Classification 66.57 77.06
WikipediaRerankingMultilingual 86.85 82.65
BelebeleRetrieval 73.89 85.18
WikipediaRetrievalMultilingual 71.90 82.94
OpusparcusPC 91.41 91.63
FinParaSTS 20.97 17.24
Avg. 66.97 72.83
Table 14: Detailed results of E5-Mistral and LUSIFER on the Finnish benchmark datasets.
Ko Datasets E5-Mistral LUSIFER
MassiveIntentClassification 70.42 69.79
MassiveScenarioClassification 75.12 75.60
KorSarcasmClassification 57.64 55.28
SIB200Classification 72.70 77.89
KorHateSpeechMLClassification 8.49 7.54
PawsXPairClassification 53.10 54.97
KLUE-TC 60.58 63.95
SIB200ClusteringS2S 31.04 46.58
Ko-StrategyQA 63.81 68.66
BelebeleRetrieval 80.09 84.69
KLUE-STS 83.48 84.17
KorSTS 79.28 78.36
STS17 80.97 80.55
Avg. 62.82 65.23
Table 15: Detailed results of E5-Mistral and LUSIFER on the Korean benchmark datasets.
Hi Datasets E5-Mistral LUSIFER
MTOPIntentClassification 68.84 79.93
SentimentAnalysisHindi 58.98 73.92
MassiveIntentClassification 64.69 71.01
MassiveScenarioClassification 69.71 75.42
SIB200Classification 68.43 75.98
TweetSentimentClassification 37.70 40.78
XNLI 65.04 60.26
IndicReviewsClusteringP2P 40.04 42.40
SIB200ClusteringS2S 27.32 45.62
WikipediaRerankingMultilingual 85.22 78.17
BelebeleRetrieval 69.73 66.76
MintakaRetrieval 18.60 21.53
MLQARetrieval 35.37 54.54
WikipediaRetrievalMultilingual 74.62 75.25
IndicCrosslingualSTS 42.30 58.97
SemRel24STS 73.14 77.34
Avg. 56.23 62.37
Table 16: Detailed results of E5-Mistral and LUSIFER on the Hindi benchmark datasets.
Bn Datasets E5-Mistral LUSIFER
BengaliDocumentClassification 50.78 48.00
BengaliHateSpeechClassification 54.67 51.43
MassiveIntentClassification 59.51 66.65
MassiveScenarioClassification 64.57 70.91
XNLIV2 63.66 60.01
IndicReviewsClusteringP2P 38.20 45.68
SIB200ClusteringS2S 23.88 43.96
WikipediaRerankingMultilingual 82.66 76.39
BelebeleRetrieval 60.17 55.77
IndicQARetrieval 56.59 68.06
WikipediaRetrievalMultilingual 71.05 72.47
IndicCrosslingualSTS 35.42 41.86
Avg. 55.10 58.43
Table 17: Detailed results of E5-Mistral and LUSIFER on the Bengali benchmark datasets.
Te Datasets E5-Mistral LUSIFER
IndicNLPNewsClassification 89.46 98.90
IndicSentimentClassification 61.53 90.63
MassiveIntentClassification 47.34 68.69
MassiveScenarioClassification 51.67 74.17
SIB200Classification 46.23 74.56
TeluguAndhraJyotiNewsClassification 67.40 76.24
IndicReviewsClusteringP2P 34.02 43.62
SIB200ClusteringS2S 10.81 42.11
BelebeleRetrieval 42.46 80.32
IndicQARetrieval 33.67 57.61
IndicCrosslingualSTS 8.36 43.76
SemRel24STS 72.83 80.99
Avg. 47.15 69.30
Table 18: Detailed results of E5-Mistral and LUSIFER on the Telugu benchmark datasets.
Sw Datasets E5-Mistral LUSIFER
AfriSentiClassification 39.67 46.47
MasakhaNEWSClassification 72.96 74.79
MassiveIntentClassification 52.84 52.79
MassiveScenarioClassification 61.09 58.59
SwahiliNewsClassification 63.95 61.56
XNLI 58.86 57.82
MasakhaNEWSClusteringP2P 34.15 36.95
MasakhaNEWSClusteringS2S 21.34 35.97
Avg. 50.61 53.12
Table 19: Detailed results of E5-Mistral and LUSIFER on the Swahili benchmark datasets.

### v2

@@ 4.2. Experimental Setup
Table 4. Number of samples used in each dataset for training. The number of negative samples is included in the total number of samples.
Baselines. We evaluate LUSIFER’s performance across the five fundamental embedding tasks on the benchmark datasets. We make comparisons with a variety of baseline models for embedding tasks which only trained/finetuned on mainly English data. Baselines include the following categories: dense retrieval models with Small Language Model (SLM) backbone: SimCSE (Gao et al., 2021a), Contriever (Izacard et al., 2022), GTE-large (Li et al., 2023), BGE-en-1.5 (Xiao et al., 2023), E5-large (Wang et al., 2024a); and dense retrieval models with Large Language Model (LLM) backbone: GTR-XXL (Ni et al., 2021b), ST5-XXL (Ni et al., 2021c), E5-Mistral (Wang et al., 2024b). Moreover, we include the following state-of-the-art multilingual embedding models which are trained on extensive multilingual data for reference: Jina-embeddings-v3 (Sturua et al., 2024), mGTE-base (Zhang et al., 2024), BGE-M3 (Chen et al., 2024), Multilingual-E5-large (Wang et al., 2024d), and UDEVER-Bloom-7B (Zhang et al., 2023).
4.3. Main Results
Table 1 presents a comprehensive evaluation of LUSIFER’s performance across 14 diverse languages, demonstrating its capabilities in multilingual representation learning. Our model achieves state-of-the-art performance in 10 out of 14 languages, with an average score of 62.63, surpassing the previous benchmark set by E5-Mistral (59.44) by 3.19 points. This improvement is particularly noteworthy given that E5-Mistral utilizes extensive proprietary synthetic data and multilingual training resources. The performance distribution across different language categories reveals LUSIFER’s robust multilingual capabilities. In high-resource languages, while maintaining competitive performance with established benchmarks, our model shows particular strength in medium and low-resource languages. The most striking improvement is observed in Telugu (te), where LUSIFER achieves a remarkable 22.15 points gain over E5-Mistral, underscoring its effectiveness in enhancing representation capabilities for traditionally underrepresented languages. When compared to existing embedding models with SLM backbones, LUSIFER demonstrates substantial improvements over models like E5-large (38.48) and BGE-en-1.5 (39.98) which are trained on English data only, thus further demonstrating the benefits of combining multilingual encoder and LLM’s English-centric for text-embedding tasks in multilingual settings.
Furthermore, LUSIFER achieves competitive performance (62.63) against state-of-the-art multilingual models such as BGE-M3 (64.80) and Multilingual-E5-large (64.54), despite these models requiring extensive multilingual training data. A key advantage of LUSIFER lies in its resource efficiency. While conventional approaches often rely heavily on extensive multilingual training data, our model achieves comparable or superior performance through its innovative alignment mechanism. This efficiency is particularly valuable in scenarios where multilingual training resources are limited or costly to obtain. The model’s ability to generalize effectively across languages while maintaining high performance demonstrates the robustness of our approach.
4.4. Cross-Lingual Evaluation
Figure 4. Performance comparison of LUSIFER and baseline models on Reranking tasks.
(a) Retrieval tasks
(b) STS tasks
Figure 5. Performance comparison of LUSIFER and baseline models on Retrieval and STS tasks.
LUSIFER demonstrates cross-lingual capabilities across multiple languages and tasks, as evidenced by the evaluation results presented in Table 2. Our model achieves a state-of-the-art average score of 57.89, surpassing the previous benchmark set by E5-Mistral (52.14) by a significant margin of 5.75 points. The performance breakdown across different datasets reveals LUSIFER’s consistent superiority. In the MLQARetrieval, LUSIFER achieves a score of 36.68, demonstrating a 5.14 improvement over E5-Mistral’s 31.54. Similarly, on the BelebeleRetrieval dataset, LUSIFER’s score of 57.81 outperforms E5-Mistral’s 54.75, showcasing its robust cross-lingual retrieval capabilities. For semantic similarity tasks, LUSIFER maintains competitive performance, achieving scores of 81.09 and 70.49 on STS17 and STS22 respectively, closely matching or slightly trailing E5-Mistral’s performance. Perhaps the most striking achievement of LUSIFER is its performance on low-resource languages, particularly evident in the IndicCrosslingual dataset. Here, LUSIFER achieves an unprecedented score of 43.40, nearly doubling E5-Mistral’s performance of 21.92. This remarkable improvement demonstrates LUSIFER’s ability to effectively transfer semantic knowledge across language boundaries, particularly benefiting languages with limited resources. These results underscore LUSIFER’s effectiveness in enhancing cross-lingual capabilities through efficient multilingual representation alignment, enabling the model to process text-embedding tasks across multiple languages effectively.
4.5. Task-Specific Performance
Figure 3, 4, 5 present the performance comparison of LUSIFER and baseline models on Classification, Clustering, Reranking, Retrieval, and STS tasks. LUSIFER consistently outperforms the baseline models across 4 out of 5 tasks, with the largest improvements observed in Clustering and Retrieval tasks, especially in the medium and low-resource languages. However, the performance of LUSIFER in the Reranking tasks is slightly worse than the strongest baseline, E5-Mistral model. This discrepancy may be attributed to the task’s complexity and the information loss in the alignment process between the multilingual encoder and the target LLM. Nevertheless, LUSIFER’s strong performance across a variety of tasks and languages highlights its ability to enhance multilingual representations without relying on explicit multilingual training data.
Baselines En Es Ru Fr Vi Fa Id Ar Fi Ko Hi Bn Te Sw Avg.
Table 5. Ablation study results of LUSIFER’s components. The table presents average metrics for each model, with the highest score for each language emphasized in bold.
4.6. Ablation Study
To comprehensively evaluate LUSIFER’s architectural design and training methodology, we conducted an extensive ablation study examining the contribution of each major component. We compared LUSIFER against several ablated variants:
  1. (1)
    Connector-Only: A simplified version using only the finetuning connector during both alignment training and representation finetuning stages, removing the complex interaction between components.
  2. (2)
    Frozen Multilingual Encoder: A variant where the multilingual encoder is freeze throughout training, with only the connector trained during alignment and both connector and target LLM trained during representation finetuning.
  3. (3)
    Alignment Only: The model trained solely with alignment training, omitting the representation finetuning stage to assess the importance of the two-stage training process.
  4. (4)
    Representation Finetuning Only: A version trained using only representation finetuning without initial alignment training.
We evaluated these variants across our standard benchmark suite, measuring performance on cross-lingual transfer, semantic similarity, and downstream task effectiveness. Table 5 presents the comprehensive results of this analysis. The full LUSIFER model achieved the highest average score of 62.63, significantly outperforming all ablated versions. Breaking down the performance impacts: (i) The Connector-Only variant (44.18) showed a 18.45 point, highlighting the importance of allowing the multilingual encoder and the target LLM to be finetuned during training. (ii) The Frozen Multilingual Encoder variant (56.74) performed better than Connector-Only but still fell short by 3.89, demonstrating the value of end-to-end training. (iii) Alignment-Only (44.45) and Representation Finetuning Only (57.28) variants both showed substantial degradation, with drops of 18.18 and 5.35 respectively, indicating the complementary nature of our two-stage training approach, especially the importance of the representation finetuning stage. Through this ablation study, we can attributed that the alignment training stage proves crucial for establishing initial cross-lingual connections, while the representation finetuning stage refines these alignments for specific downstream tasks.
4.7. Model Representation Visualization
(a) E5-Mistral (language)
(b) LUSIFER (language)
Figure 6. t-SNE representation of 200 randomly samples from the SIB200 dataset. The points are colored by the languages.
Figure 6 shows 2D scatter plots of representations from different models for 200 randomly sampled examples from the SIB200 dataset, visualized using t-SNE. The points are colored by the language of the samples. The t-SNE representation of E5-Mistral demonstrates a clearer separation between languages, with distinct clusters for each language. In contrast, the visualization of LUSIFER presents a more mixed distribution of languages, with overlapping clusters across different languages. This observation provides insights into LUSIFER’s lingual-agnostic capabilities, highlighting the model’s ability to bridge the gaps between representation spaces of different languages. These results suggest that LUSIFER’s alignment strategy enables the model to comprehend semantics across multiple languages effectively, facilitating zero-shot multilingual transfer. Overall, our experiments confirm the advantages of the representation alignment strategies in LUSIFER in effectively enabling zero-shot multilingual transfer for LLM-based embedding methods.
5. Conclusion
In this work, we propose LUSIFER, a novel framework that enables effective multilingual representation without explicit multilingual training data. LUSIFER aligns a multilingual encoder with a target English-centric LLM through a minimal set of trainable parameters, facilitating zero-shot multilingual transfer. Our experimental results demonstrate that LUSIFER achieves state-of-the-art performance across diverse languages and tasks, outperforming existing baseline models. Moreover, LUSIFER significantly enhances cross-lingual capabilities, enabling the model to process text-embedding tasks across multiple languages effectively.
Our work provides a promising direction for enhancing multilingual representation capabilities in English-centric embedding models, enabling global applicability without requiring extensive multilingual training data. The alignment strategies employed in LUSIFER ensure that the model can comprehend and process semantic information across different languages, making it a versatile tool for various multilingual applications. Furthermore, the ability of LUSIFER to perform zero-shot multilingual transfer opens up new possibilities for natural language processing tasks in low-resource languages, where obtaining large-scale annotated data is often challenging. By leveraging the strengths of both multilingual encoders and English-centric LLMs, LUSIFER bridges the gap between languages, fostering better multilingual representation learning.
In future work, we plan to explore additional alignment strategies and further investigate the impact of LUSIFER’s components on multilingual representation quality. We also aim to extend our framework to support different modalities such as images and audio and evaluate its performance on a wider range of tasks. Additionally, we will explore the integration of LUSIFER with other state-of-the-art models to further enhance its capabilities. Moreover, we anticipate that LUSIFER will facilitate broader applications of LLM embeddings in downstream tasks, ranging from deep context understanding requirements like sentiment analysis (Gupta et al., 2024) to text style comprehension tasks such as authorship attribution (Rivera-Soto et al., 2021; Man and Nguyen, 2024), thereby contributing to the advancement of natural language processing and information retrieval fields.
Acknowledgements
This research has been supported by the NSF grant # 2239570. This research is also supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract 2022-22072200003. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of ODNI, IARPA, or the U.S. Government. The U.S. Government is authorized to reproduce and distribute reprints for governmental purposes, notwithstanding any copyright annotation therein.